The point P on the unit circle that corresponds to a real number t is [5/7, -2sqrt6/7] Find csc (t)
Answer by lwsshak3(6494)

The point P on the unit circle that corresponds to a real number t is [5/7, -2sqrt6/7] Find csc (t)
**
In a unit circle, the x-coordinate=cos(t) and y-coordinate=sin(t)
sin(t)= -2√6/7
csc(t)=1/sin(t)=1/(-2√6/7)=7/(-2√6)
rationalize denominator, multiply top and bottom by √6
csc(t)=7/(-2√6)=-(7√6)/12